FINE-STRUCTURE PREDISSOCIATION OF ULTRACOLD PHOTOASSOCIATED ^{39}K_{2} MOLECULES

\underline{HE WANG}, P. L. GOULD and W. C. STWALLEY, Department of Physics, University of Connecticut, Storrs, CT 06269.

We report the first observation of fragmentation spectra of the fine-structure predissociated long-range 1_{g} and 0_{u}^{+} state of ultracold (T - 350μK) ^{39}K_{2} molecules. The 1_{g} state predissociates via the curve crossing at long range with the repulsive 0_{g}^{+} state. The 0_{u}^{+} state, however, predissociates predominantly through molecular fine-structure mixing at short range with the b^{3}Π_{1u} component. Our technique of fragmentation spectroscopy is a sensitive and selective tool for studies of ultracold photoassociation and fine-structure-changing collisions.